OpenStudy (anonymous):

Let f(x)=x^5+3x-2. Then the inverse of f(x) (2) equals a) 1/83 b)1/8 c) 1 d)8 e)83 Please provide a detailed explanation

OpenStudy (saifoo.khan):

Do you know how to find the inverse of a function?

OpenStudy (anonymous):

Yes, I do but I keep getting the answer of a) 1/83 and my teacher said the correct answer is e)83

OpenStudy (saifoo.khan):

Can you show me your working?

OpenStudy (anonymous):

x= y^5 +3y-2 * take derivative x=5y^4 + 3 1/5(2)^4 + 3= 1/83 because after you take the derivative of your inverse I was told you put in under 1

OpenStudy (saifoo.khan):

Umm. I don't get this working.

OpenStudy (saifoo.khan):

The derivative will be 5y^4 + 3.

OpenStudy (anonymous):

okay

OpenStudy (saifoo.khan):

But idk how to find inverse using derivates.

OpenStudy (saifoo.khan):

derivatives*